# Exploring Test-Time Adaptation for Object Detection in Continually Changing Environments

### v3

@@ Abstract.
Real-world application models are commonly deployed in dynamic environments, where the target domain distribution undergoes temporal changes. Continual Test-Time Adaptation (CTTA) has recently emerged as a promising technique to gradually adapt a source-trained model to continually changing target domains. Despite recent advancements in addressing CTTA, two critical issues remain: 1) Fixed thresholds for pseudo-labeling in existing methodologies generate low-quality pseudo-labels, as model confidence varies across categories and domains; 2) Stochastic parameter restoration methods for mitigating catastrophic forgetting fail to effectively preserve critical information due to their intrinsic randomness. To tackle these challenges for detection models in CTTA scenarios, we present CTAOD, featuring three core components. Firstly, the object-level contrastive learning module extracts object-level features for contrastive learning to refine the feature representation in the target domain. Secondly, the adaptive monitoring module dynamically skips unnecessary adaptation and updates the category-specific threshold based on predicted confidence scores to enable efficiency and improve the quality of pseudo-labels. Lastly, the data-driven stochastic restoration mechanism selectively reset inactive parameters with higher possibilities, ensuring the retention of essential knowledge. We demonstrate the effectiveness of CTAOD on four CTTA object detection tasks, where CTAOD outperforms existing methods, especially achieving a 3.2 mAP improvement and a 20% increase in efficiency on the Cityscapes-to-Cityscapes-C CTTA task. The code will be released.
@@ 1. Introduction
Figure 1. Motivation. The source-trained model adapts to dynamic environments without source data. (1) AM: The bar chart shows mean model scores varying across categories and domains. Unlike existing methods that continually adapt to all test data using fixed thresholds, CTAOD dynamically skips unnecessary adaptations and updates the category-specific threshold. (2) DSR: The bottom illustrates the neuron reset process, where darker purple indicates higher activity, and blue represents source knowledge. While existing methods restore parameters randomly, CTAOD resets inactive parameters with higher possibilities.
@@
Deep learning models have demonstrated immense potential across various data modalities such as image (He et al., 2016; Dosovitskiy et al., 2020), audio (Aytar et al., 2016), and language (Devlin et al., 2018). However, these models experience pronounced degradation in performance when confronted with training data (i.e., source domain) and testing data (i.e., target domain) originating from disparate distributions. This phenomenon, commonly referred to as distribution shifts, poses a significant challenge (Geirhos et al., 2018; Hendrycks and Dietterich, 2019).
In such a scenario, unsupervised domain adaptation (UDA) becomes crucial, which typically involves aligning the source and target data distributions, thereby mitigating the impact of distribution shifts (Ganin et al., 2016; Tzeng et al., 2017). Still, UDA falls short by necessitating access to source data, which is often inaccessible due to privacy constraints or data transmission barriers (Huang et al., 2021).
@@
This limitation catalyzes the exploration of Test-Time Adaptation (TTA) where the source-trained model directly adapts toward unlabeled test samples encountered during evaluation in an online manner, without the reliance on the source data (Wang et al., 2020). Nonetheless, the TTA methods, which assume a static target domain, face a more challenging and realistic problem, as real-world systems work in non-stationary environments. For example, a vehicle may encounter various continuous environmental changes such as fog, night, rain, and snow during its journey. Existing TTA methods are vulnerable to catastrophic forgetting of source knowledge and error accumulation when adaptation faces more than one distribution shift (Wang et al., 2022).
@@
Recently, Wang et al. (2022) introduce CoTTA by applying stochastic parameters restoration to mitigate catastrophic forgetting in Continual Test-Time Adaptation (CTTA) scenarios, where the model is continually adapted to sequences of target domains. Although randomly resetting parameters partially helps alleviate the forgetting of source knowledge (Wang et al., 2022; Zhu et al., 2023), its randomness may also contribute to losing crucial knowledge specific to the current domain (refer to Figure 1). Additionally, existing CTTA methods normally utilize pseudo-labeling through a fixed threshold for self-supervision (Wang et al., 2022; Döbler et al., 2023). Given that model confidence may vary across categories and domains (refer to Figure 1), employing a uniformly fixed threshold could exclude high-quality pseudo-labels while incorporating incorrect ones. Furthermore, these methods adapt to every new incoming data, resulting in computational inefficiency and potential performance degradation. Low-quality pseudo-labels and unstable adaptation to noisy data lead to error accumulation, as they provide negative feedback to the model.
@@
To tackle these challenges, we introduce CTAOD (Continual Test-time Adaption for Object Detection). Aligning with previous CTTA works for robust adaptation (Wang et al., 2022), CTAOD is constructed upon the mean-teacher framework (Tarvainen and Valpola, 2017) where the teacher model is an exponential moving average of the student model. Particularly, CTAOD comprises Object-level Contrastive Learning (OCL), Adaptive Monitoring (AM), and Data-driven Stochastic Restoration (DSR).
More specifically, OCL extracts features based on proposals generated by the Region Proposal Network (RPN), which provides multiple cropped views around the object at different locations and scales. Subsequently, Contrastive Learning (CL) loss is applied on RPN cropped views to guide the model to encourage similar object instances to remain close while pushing dissimilar ones apart. The OCL is well integrated into the mean-teacher paradigm as a drop-in enhancement to acquire more fine-grained and localized feature representation. Furthermore, to make adaptation more efficient and stable and to improve the quality of pseudo-labels, i.e., addressing Challenge 1, we design an AM module to decide whether to pause or resume the adaptation and dynamically adjust the category-specific thresholds, based on the predicted confidence scores. The dynamic nature of the AM method makes it better suited to address the effects of continuously changing distributions. Finally, the DSR mechanism is proposed to reset inactive parameters with a higher possibility than active ones by employing the gradients of parameters as weights for a random mask matrix. DSR not only helps prevent forgetting but preserves important information, i.e., addressing Challenge 2. On the other hand, its randomness also allows potentially falsely activated parameters to be reset, thereby making the model more stable.
@@
  • This study introduces CTAOD, which pioneers exploring CTTA for detection models. Specifically, we propose to leverage object-level features for contrastive learning to refine feature representation in CTTA object detection.
  • To address the two challenges in CTTA, our proposed CTAOD enables dynamic skipping and category-specific threshold updates based on predicted scores and reset the inactive parameter with higher possibilities.
  • Empirical experiments demonstrate that CTAOD surpasses existing methods and facilitates efficient short-term and long-term adaptation in dynamic environments.
@@ 2.1. Source-Free Domain Adaptation
UDA tackles the inter-domain divergence by aligning the distributions of source and target data (Ganin et al., 2016; Luo et al., 2020). Despite its effectiveness, the limitation of UDA lies in its requirement for access to the source domain data, which often raises concerns regarding data privacy and transmission efficiency. As a result, Source-Free Domain Adaptation (SFDA) received extensive research attention in object detection, where the source-trained detector is re-trained on the target training data (VS et al., 2023b, a; Li et al., 2021) before evaluation. For instance, MemCLR (VS et al., 2023a) employs a cross-attention-based memory bank for CL, while IRG (VS et al., 2023a) utilizes the object relations with instance relation graph network to explore the SFDA setting for object detection without the source data. However, the standard SFDA setting requires prior knowledge of the target domain, which is impractical in most real-world applications.
@@ 2.2. Test-Time Adaptation
TTA adapts the source-trained model to the target test data during inference time without access to the source data. Since both TTA and SFDA involve adapting the source-trained model to the unlabeled target data without utilizing source data, some works also refer to TTA as SFDA (Wang et al., 2022; Brahma and Rai, 2023). In this paper, we distinguish between TTA and SFDA based on evaluation protocol, although they can be transformed into each other in experiments. Furthermore, TTA methods improve the model performance under distribution shift commonly through pseudo-labeling (Sinha et al., 2023; Iwasawa and Matsuo, 2021), batchnorm statistics updating (Hu et al., 2021; You et al., 2021), or entropy regularization (Wang et al., 2020; Iwasawa and Matsuo, 2021) during testing. For example, Tent (Wang et al., 2020) updates the batchnorm parameters with entropy minimization and demands a large batch size for optimization during test-time adaptation, which is unsuitable for real-time detection model deployment where images are processed sequentially. TTA methods assume a static target domain where the target data come from a single domain. However, the target distribution may exhibit a continual shift over time in practical scenarios.
@@ 2.3. Continual Test-time Adaptation
Recently, Wang et al. (2022) introduce CoTTA, marking the first work tailored to the demands of CTTA by adapting a pre-trained model to sequences of domains without source data. Subsequently, research efforts have been dedicated to exploring CTTA, primarily focused on classification (Döbler et al., 2023; Brahma and Rai, 2023; gan2023decorate) and segmentation (Song et al., 2023; Ni et al., 2023; Liu et al., 2023) tasks. For instance, similar to CoTTA, Zhu et al. (2023) apply a stochastic reset mechanism to prevent forgetting in the CTTA medical segmentation task. In contrast, PETAL (Brahma and Rai, 2023) utilizes the Fisher Information Matrix (FIM) as a metric of parameter importance to reset only the most irrelevant parameters. Nevertheless, these two methods are sub-optimal, as the randomness might lead to losing essential information, and pure data-driven restoration may retain false active parameters resulting from noise.
Moreover, the mean-teacher framework (Tarvainen and Valpola, 2017) serves as a base architecture for most CTTA works (Wang et al., 2022; Gong et al., 2022), where the teacher generates pseudo-labels via a fixed threshold to supervise the student. Nonetheless, these methods suffer from low-quality pseudo-labels with a uniform threshold since the model confidence varies across categories and domains. Furthermore, Wang et al. (2024) design a dynamic thresholding technique to update the threshold in a batch for classification tasks, requiring a large batch size. However, this approach is unsuitable for object detection where smaller batch sizes are preferred for computational efficiency. To improve efficiency, yoo2024and presents a CTTA approach for object detection to update specific parts of the model in specific adaptation, which utilizes source statistics. Although it effectively improved efficiency, its performance is limited. Additionally, Gan et al. (2023) presents a cloud-device collaborative continual adaptation paradigm in object detection by aligning source and target distribution. Nevertheless, the reliance on source data in both methods restricts their practical applicability. Therefore, a gap still exists in exploring the CTTA in object detection to improve the quality of pseudo-label (Challenge 1) and effectively reset noisy neurons (Challenge 2), without relying on any source data.
@@ 3.1.1. Problem Statement
Figure 2. Overview. CTAOD follows the mean-teacher framework, featuring: (1) The OCL module compares region of interest features extracted from feature maps of both networks based on teacher proposals for contrastive learning; (2) The AM module dynamically skips unnecessary adaptation and adjusts category-specific thresholds based on prediction scores; (3) The DSR reset the inactive parameter with higher possibilities based on the FIM.
@@ 3.1.2. Mean-Teacher Framework
In alignment with previous CTTA works (Wang et al., 2022; Gong et al., 2022), we build our method based on the mean-teacher framework (Tarvainen and Valpola, 2017). Specifically, both networks are first initialized with the source-trained model. The teacher model produces the pseudo-labels for the unlabeled target data to supervise the student model. While the parameters of the student model are optimized via gradient descent, the teacher model is updated following an Exponential Moving Average (EMA) strategy based on the student model. Formally, this process can be expressed as follows:
@@
where $x_{t}$ and $\hat{y}_{t}$ denote the unlabeled target data and corresponding pseudo-labels. $\theta^{std}_{t}$ and $\theta^{tch}_{t}$ symbolize the parameters of student and teacher networks at time step $t$. Moreover, the supervision loss $\mathcal{L}_{pl}$ in Faster-RCNN (Ren et al., 2015), which consists of the RPN loss $\mathcal{L}_{rpn}$ and RCNN loss $\mathcal{L}_{rcnn}$, is utilized for pseudo-labeling. Additionally, $\gamma$ represents the student’s learning rate, and the EMA rate is denoted by $\alpha$. Hence, the teacher can be regarded as an ensemble of historical students, providing stable supervision. However, this framework encounters error accumulation and catastrophic forgetting in dynamic environments. Therefore, we design CTAOD to fit in the CTTA setting.
@@ 3.2.1. Overview
Figure 2 presents a concise overview of CTAOD, consisting of Object-level Contrastive Learning (OCL), Adaptive Monitoring (AM), and Data-driven Stochastic Restoration (DSR). Inspired by UDA object detection methods (Li et al., 2022; Liu et al., 2021), we adopt the Weak-Strong augmentation to enable the teacher model to generate reliable pseudo-labels without being affected by heavy augmentation. Specifically, the teacher model receives input with weak augmentations while we input student networks with strong enhanced images. The overall loss of the student model for gradient descent is defined as:
@@
where $\mathcal{L}_{cl}(x_{t})$ denotes the contrastive learning (CL) loss, to be elaborated later. The $\mathcal{L}_{kl}$ represents the Kullback-Leibler (KL) Divergence (Kullback and Leibler, 1951) loss, used to quantify the distinction between two probability distributions. The parameters $\lambda$ and $\mu$ are corresponding weights of the losses. The KL loss is defined as:

### v4

@@ Abstract.
Real-world application models are commonly deployed in dynamic environments, where the target domain distribution undergoes temporal changes. Continual Test-Time Adaptation (CTTA) has recently emerged as a promising technique to gradually adapt a source-trained model to continually changing target domains. Despite recent advancements in addressing CTTA, two critical issues remain: 1) Fixed thresholds for pseudo-labeling in existing methodologies lead to low-quality pseudo-labels, as model confidence varies across categories and domains; 2) Stochastic parameter restoration methods for mitigating catastrophic forgetting fail to preserve critical information effectively, due to their intrinsic randomness. To tackle these challenges for detection models in CTTA scenarios, we present AMROD, featuring three core components. Firstly, the object-level contrastive learning module extracts object-level features for contrastive learning to refine the feature representation in the target domain. Secondly, the adaptive monitoring module dynamically skips unnecessary adaptation and updates the category-specific threshold based on predicted confidence scores to enable efficiency and improve the quality of pseudo-labels. Lastly, the adaptive randomized restoration mechanism selectively reset inactive parameters with higher possibilities, ensuring the retention of essential knowledge. We demonstrate the effectiveness of AMROD on four CTTA object detection tasks, where AMROD outperforms existing methods, especially achieving a 3.2 mAP improvement and a 20% increase in efficiency on the Cityscapes-to-Cityscapes-C CTTA task. The code will be released.
@@ 1. Introduction
Figure 1. The Motivation of our proposed method. The source-trained model adapts to continually changing target domains without source data. (1) AM: The bar chart shows mean model scores varying across categories and domains. Unlike existing methods that continually adapt to all test data using fixed thresholds, AMROD dynamically skips unnecessary adaptations and updates the category-specific threshold. (2) ARR: The bottom illustrates the neuron reset process, where darker purple indicates higher activity, and blue represents source knowledge. While existing methods restore parameters randomly, AMROR resets inactive parameters with higher possibilities.
@@
Deep learning models have demonstrated immense potential across various data modalities such as image (He et al., 2016; Dosovitskiy et al., 2020), audio (Aytar et al., 2016), and language (Devlin et al., 2018). However, these models experience pronounced degradation in performance when confronted with training data (i.e., source domain) and testing data (i.e., target domain) originating from disparate distributions. This phenomenon, commonly referred to as distribution shifts, poses a significant challenge (Geirhos et al., 2018; Hendrycks and Dietterich, 2019). In such a scenario, unsupervised domain adaptation (UDA) becomes crucial, which typically involves aligning the source and target data distributions, thereby mitigating the impact of distribution shifts (Ganin et al., 2016; Tzeng et al., 2017). Still, UDA falls short by necessitating access to source data, which is often inaccessible due to privacy constraints or data transmission barriers (Huang et al., 2021).
@@
This limitation catalyzes the exploration of Test-Time Adaptation (TTA) where the source-trained model directly adapts toward unlabeled test samples encountered during evaluation in an online manner, without the reliance on the source data (Mummadi et al., 2021; Sinha et al., 2023; You et al., 2021; Wang et al., 2020). Nonetheless, the TTA methods, which assume a static target domain, face a more challenging and realistic problem, as real-world systems work in non-stationary environments. For example, a vehicle may encounter various continuous environmental changes such as fog, night, rain, and snow during its journey. Existing TTA methods are vulnerable to catastrophic forgetting of source knowledge and error accumulation when adaptation faces more than one distribution shift (Wang et al., 2022; Döbler et al., 2023).
@@
Recently, Wang et al. (2022) introduce CoTTA by applying stochastic parameters restoration to mitigate catastrophic forgetting in Continual Test-Time Adaptation (CTTA) scenarios, where the model is continually adapted to sequences of target domains. Although randomly resetting parameters partially helps alleviate the forgetting of source knowledge (Wang et al., 2022; Zhu et al., 2023), its randomness may also contribute to losing crucial knowledge specific to the current domain (refer to Figure 1). Additionally, existing CTTA methods normally utilize pseudo-labeling through a fixed threshold for self-supervision (Wang et al., 2022; Döbler et al., 2023; Sójka et al., 2023). Given that model confidence may vary across categories and domains (refer to Figure 1), employing a uniformly fixed threshold could exclude high-quality pseudo-labels while incorporating incorrect ones. Furthermore, these methods adapt to every new incoming data, resulting in computational inefficiency and potential performance degradation. Low-quality pseudo-labels and unstable adaptation to noisy data lead to error accumulation, as they provide negative feedback to the model.
@@
To tackle these challenges, we introduce AMROD (Adaptive Monitoring and Restoration for Object Detection). Aligning with previous CTTA works for robust adaptation (Wang et al., 2022; Gong et al., 2022; Döbler et al., 2023; Niloy et al., 2024), AMROR is constructed upon the mean-teacher framework (Tarvainen and Valpola, 2017) where the teacher model is an exponential moving average of the student model. Particularly, AMROR comprises Object-level Contrastive Learning (OCL), Adaptive Monitoring (AM), and Adaptive Randomized Restoration (ARR) modules.
More specifically, OCL extracts object-level features based on proposals generated by the Region Proposal Network (RPN), which provides multiple cropped views around the object at different locations and scales. Subsequently, Contrastive Learning (CL) loss is applied on RPN cropped views to guide the model to encourage similar object instances to remain close while pushing dissimilar ones apart. The OCL is well integrated into the mean-teacher paradigm as a drop-in enhancement to acquire more fine-grained and localized feature representation. Furthermore, to make adaptation more efficient and stable and to improve the quality of pseudo-labels, i.e., addressing Challenge 1, we design an AM module to decide whether to pause or resume the adaptation and dynamically adjust the category-specific thresholds, based on the mean predicted confidence scores. The dynamic nature of the AM method makes it better suited to address the effects of continuously changing distributions. Finally, the ARR mechanism is proposed to reset inactive parameters with a higher possibility than active ones by employing the gradients of parameters as weights for a random mask matrix. ARR not only helps prevent forgetting but preserves important information, i.e., addressing Challenge 2. On the other hand, its randomness also allows certain falsely activated parameters to be reset as well, thereby resulting in greater stability for adaptation.
@@
  • This study introduces AMROR, which pioneers exploring CTTA for detection models. Specifically, we propose to leverage object-level features for contrastive learning to refine feature representation in CTTA object detection.
  • To address the two challenges in CTTA, our proposed AMROR enables dynamic skipping and category-specific threshold updates based on the mean predicted scores and reset the inactive parameter with higher possibilities.
  • Empirical experiments demonstrate that AMROR surpasses existing methods and facilitates efficient short-term and long-term adaptation in dynamic environments.
@@ 2.1. Source-Free Domain Adaptation
UDA tackles the inter-domain divergence by aligning the distributions of source and target data (Ganin et al., 2016; Tzeng et al., 2017; Saito et al., 2018, 2019; Li et al., 2022a; Cao et al., 2023). Despite its effectiveness, the limitation of UDA lies in its requirement for access to the source domain data, which often raises concerns regarding data privacy and transmission efficiency. As a result, Source-Free Domain Adaptation (SFDA) received extensive research attention in object detection, where the source-trained detector is adapted to the target training data without any source data (VS et al., 2023b, a; Li et al., 2022b, 2021; Chen et al., 2023) before evaluation. For instance, MemCLR (VS et al., 2023a) employs a cross-attention-based memory bank with CL for source-free detectors, while IRG (VS et al., 2023a) utilizes the object relations with instance relation graph network to explore the SFDA setting for object detection. However, the standard SFDA setting requires prior knowledge of the target domain, which is impractical in most real-world applications.
@@ 2.2. Test-Time Adaptation
TTA adapts the source-trained model to the target test data during inference time without access to the source data. Since both TTA and SFDA involve adapting the source-trained model to the unlabeled target data without utilizing source data, some works also refer to TTA as SFDA (Wang et al., 2022; Brahma and Rai, 2023). In this paper, we distinguish between TTA and SFDA based on evaluation protocol, although they can be transformed into each other in experiments. Furthermore, TTA methods improve the model performance under distribution shift commonly through pseudo-labeling (Mummadi et al., 2021; Sinha et al., 2023; Iwasawa and Matsuo, 2021; Sun et al., 2020), batchnorm statistics updating (Hu et al., 2021; You et al., 2021), or entropy regularization (Wang et al., 2020; Zhou and Levine, 2021; Iwasawa and Matsuo, 2021; Niu et al., 2022) during testing. For example, Tent (Wang et al., 2020) updates the batchnorm parameters with entropy minimization and demands a large batch size for optimization during test-time adaptation, which is unsuitable for real-time detection model deployment where images are processed sequentially. The above approaches assume a static target domain where the target data come from a single domain. However, in numerous practical scenarios, the distribution of target domain may exhibit a continual shift over time.
@@ 2.3. Continual Test-time Adaptation
Early works consider adaptation to continually changing domains by aligning the source and target data distribution (Hoffman et al., 2014; Wulfmeier et al., 2018) and the reliance on source data limits their applicability. Recently, Wang et al. (2022) introduce CoTTA, marking the first work tailored to the demands of CTTA by adapting a pre-trained model to sequences of domains without source data. Subsequently, research efforts have been dedicated to exploring CTTA, primarily focused on classification (Döbler et al., 2023, 2023; Niloy et al., 2024; Brahma and Rai, 2023; Gan et al., 2023a) and segmentation (Song et al., 2023; Ni et al., 2023; Liu et al., 2023; Niloy et al., 2024; Zhu et al., 2023) tasks. For instance, similar to CoTTA, Zhu et al. (2023) apply a stochastic reset mechanism to prevent forgetting in the CTTA medical segmentation task. In contrast, PETAL (Brahma and Rai, 2023) utilizes the Fisher Information Matrix (FIM) as a metric of parameter importance to reset only the most irrelevant parameters across all layers. Nevertheless, these two methods are sub-optimal, as the randomness might lead to losing essential information, and pure data-driven restoration may retain false active parameters resulting from noise.
Moreover, the mean-teacher framework (Tarvainen and Valpola, 2017) serves as a base architecture for most CTTA works (Wang et al., 2022; Gong et al., 2022; Döbler et al., 2023; Niloy et al., 2024), where the teacher model generates pseudo-labels via a fixed threshold to supervise the training of the student model. Nonetheless, these methods suffer from low-quality pseudo-labels with a uniform threshold since the model confidence varies across categories and domains. Furthermore, Wang et al. (2024) design a dynamic thresholding technique to update the threshold in a batch for classification tasks, requiring a large batch size. However, this approach is unsuitable for object detection where smaller batch sizes are preferred for computational efficiency. To improve efficiency, Yoo et al. (2024) presents a CTTA approach for object detection to update specific parts of the model in specific adaptation, which utilizes source statistics. Although it effectively improved efficiency, its performance is limited. Additionally, Gan et al. (2023b) presents a cloud-device collaborative continual adaptation paradigm in object detection by aligning source and target distribution. Nevertheless, the reliance on source data in both methods restricts their practical applicability. Therefore, a gap still exists in exploring the CTTA in object detection to improve the quality of pseudo-label (Challenge 1) and effectively reset noisy neurons (Challenge 2), without relying on the source data.
@@ 3.1.1. Problem Statement
Figure 2. The overview of the proposed method. AMROD follows the mean-teacher framework, featuring: (1) The OCL module compares region of interest features extracted from feature maps of both networks based on teacher proposals for contrastive learning; (2) The AM module dynamically skips unnecessary adaptation and adjusts category-specific thresholds based on the mean prediction scores; (3) The ARR strategy reset the inactive parameter with higher possibilities based on the FIM.
@@ 3.1.2. Mean-Teacher Framework
Following previous CTTA works (Wang et al., 2022; Gong et al., 2022), we build AMROD based on the mean-teacher framework (Tarvainen and Valpola, 2017). Specifically, both networks are first initialized with the source-trained model. The teacher model produces the pseudo-labels for the unlabeled target data to supervise the student model. While the parameters of the student are optimized via gradient descent, the teacher is updated following an Exponential Moving Average (EMA) strategy based on the student. Formally, this process can be expressed as follows:
@@
where $x_{t}$ and $\hat{y}_{t}$ denote the unlabeled target data and corresponding pseudo-labels. $\theta^{std}_{t}$ and $\theta^{tch}_{t}$ symbolize the parameters of student and teacher networks at time step $t$. Moreover, the supervision loss $\mathcal{L}_{pl}$ in Faster-RCNN (Ren et al., 2015), which consists of the RPN loss $\mathcal{L}_{rpn}$ and RCNN loss $\mathcal{L}_{rcnn}$, is utilized for pseudo-labeling. Additionally, $\gamma$ represents the student’s learning rate, and the EMA rate is denoted by $\alpha$. Hence, the teacher can be regarded as an ensemble of historical students, providing stable supervision. However, this framework encounters error accumulation and catastrophic forgetting in dynamic environments. Therefore, we design AMROD to fit in the CTTA setting.
@@ 3.2.1. Overview
Figure 2 presents a concise overview of AMROD, consisting of Object-level Contrastive Learning (OCL), Adaptive Monitoring (AM), and Adaptive Randomized Restoration (ARR). Inspired by UDA object detection methods (Li et al., 2022a; Liu et al., 2021), we adopt the Weak-Strong augmentation to enable the teacher model to generate reliable pseudo-labels without being affected by heavy augmentation. Specifically, the teacher receives input with weak augmentations while we input student networks with strong enhanced images. The overall loss of the student for gradient descent is defined as:
@@
where $\mathcal{L}_{cl}(x_{t})$ denotes the contrastive learning (CL) loss, to be elaborated later. The $\mathcal{L}_{kl}$ represents the Kullback-Leibler (KL) Divergence (Kullback and Leibler, 1951) loss, used to quantify the distinction between two probability distributions. The parameters $\lambda$ and $\mu$ are corresponding weights. KL loss is defined as: